OpenStudy (anonymous):

simplyfy (x-9)^3/5=8 how do you solve the power to 3/5?

OpenStudy (anonymous):

(x-9)^(3/5) = 8 \[\sqrt[5]{(x-9)^{3}} = 8\] (x-9)³=8^5=32768 x-9=cbrt(32768)=32 x=41

OpenStudy (anonymous):

what is cbrt?

OpenStudy (anonymous):

cube root, the inverse of the power to 3.

OpenStudy (anonymous):

x-9=cbrt(32768)=32 i dont understand this part

OpenStudy (anonymous):

please help!